# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 2.2. Analogue of the Deligne-Ihara conjecture and its consequence
Note that the submodule $\Pi_{1,1}(m+3)\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ corresponds to the ideal $J^{m+1}$, where $J$ denotes the augmentation ideal of $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ [Iha86, (19) on page 67]. It follows that the power series $T_{1}\alpha_{1,1}(\sigma)$ is contained in $J^{m+1}$, and hence $\alpha_{1,1}(\sigma)\in J^{m}$. Hence every coefficient of $\alpha_{1,1}(\sigma)$ of a monomial of total degree less than $m$ vanishes. By observing Theorem 2.3, it follows that $\kappa_{\boldsymbol{n}}(\sigma)$ vanishes for every $\boldsymbol{n}\in I$ such that $\lvert\boldsymbol{n}\rvert\leq m-2$, as desired. ∎
The Deligne-Ihara conjecture [Sha02] states that a $\mathbb{Q}_{p}$-graded Lie algebra, defined in a similar manner to $\mathfrak{g}\otimes\mathbb{Q}_{p}$ using the pro-$p$ outer Galois representation of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$, is a free graded Lie algebra in each odd degree $m\geq 3$. We formulate an analogue of this conjecture, as follows:
Conjecture 2.8.
For every $\boldsymbol{m}\in I$, choose $\sigma_{\boldsymbol{m}}$ such that
  1. (1)
    The element $\sigma_{\boldsymbol{m}}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component of $\mathfrak{g}_{\boldsymbol{\lvert m\rvert}}$, and
  2. (2)
    The image $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})\in\mathbb{Z}_{p}(\boldsymbol{m})$ under $\kappa_{\boldsymbol{m}}$ is nonzero, and generates $\kappa_{\boldsymbol{m}}(\mathfrak{g}_{|\boldsymbol{m}|}).$
Then the graded Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$ is freely generated by $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$.
Remark 2.9.
(1) Using the theory of weighted completions of profinite groups developed by Hain-Matsumoto [HM03], one can prove the generation part of Conjecture 2.8, assuming that $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite for every $\boldsymbol{m}\in I$. At the writing of this paper, however, we do not know how to prove the freeness portion of the conjecture. We hope to discuss the structure of $\mathfrak{g}\otimes\mathbb{Q}_{p}$ in future research.
(2) For the second assumption of the Conjecture 2.8, it is enough to only assume $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})\neq 0$, since we concentrated on the structure of the $\mathbb{Q}_{p}$-graded Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$. However, the latter condition becomes essential when considering the structure of the $\mathbb{Z}_{p}$-graded Lie algebra $\mathfrak{g}$. Our proof of Theorem A shows that, under the assumption of Theorem A, one can choose the set $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying two assumptions of Conjecture 2.8 so that it even freely generates $\mathfrak{g}$.
We turn our attention to field-theoretic properties of the fixed field of the kernel $\ker(\rho_{X,p})$ of the pro-$p$ outer Galois representation of the once-punctured elliptic curve $X=E-O$.
@@
The fixed field $\bar{K}^{\ker(\rho_{X,p})}$ is a pro-$p$ extension of $K(E[p])$ unramified outside $p$.
@@
It suffices to show that $X\times_{K}K(E[p])$ has good reduction everywhere outside $p$, by virtue of specialization homomorphisms of étale fundamental groups and the fact that $\ker\left(\mathrm{Out}(\Pi_{1,1})\to\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)\right)$ is pro-$p$. By the Néron-Ogg-Shafarevich criterion, it suffices to show that every inertia subgroup $I$ associated to a prime of $K(E[p])$ outside $p$ acts trivially on $T_{p}(E)$.
Since $E$ has complex multiplication, it has everywhere potentially good reduction. Hence the image of $I$ in $\mathrm{Aut}(T_{p}(E))$ is finite. Moreover, the image of $I$ is contained in the subgroup $\ker(\mathrm{GL}(T_{p}(E))\to\mathrm{GL}(E[p]))$, which is a torsion-free pro-$p$ group for every $p\geq 5$. Hence the image of $I$ is trivial. ∎
@@
The field $\bar{K}^{\ker(\rho_{X,p})}$ can be written as a compositum of $K(E[p])$ and a subfield $\Omega^{\ast}\subset\Omega$, in such a way that $\mathrm{Gal}(\bar{K}^{\ker(\rho_{X,p})}/K(p))$ splits into the direct product of $\mathrm{Gal}(K(E[p])/K(p))$ and $\mathrm{Gal}(\Omega^{\ast}/K(p))$.
@@
$1\to\rho_{X,p}(G_{K(E[p])})\to\rho_{X,p}(G_{K(p)})\to\mathrm{Gal}(K(E[p])/K(p))\to 1$
Since $\rho_{X,p}(G_{K(E[p])})$ is a pro-$p$ group and $\mathrm{Gal}(K(E[p])/K(p))$ is a finite prime-to-$p$ group (here we use $p\geq 5$), this sequence splits by the Schur-Zassenhaus theorem [Ser97, 5.9, Corollary 1]. Let $t\colon\mathrm{Gal}(K(E[p])/K(p))\to\rho_{X,p}(G_{K(p)})$ be an arbitrary section.
There is a natural homomorphism $\mathrm{Aut}_{K}(X)\to\mathrm{Out}(\Pi_{1,1})$x, and the image of $\mathrm{Aut}_{K}(X)$ centralizes $\rho_{X,p}(G_{K})$.
Through the $G_{K}$-action on $E[p]$, the Galois group $\mathrm{Gal}(K(E[p])/K(p))$ is identified with a subgroup of the image of $\mathrm{Aut}_{K}(X)$ in $\mathrm{Aut}_{O_{K}}(E[p])=(O_{K}/p)^{\times}$. Moreover $\mathrm{Aut}_{K}(X)=O_{K}^{\times}$ injects into $(O_{K}/p)^{\times}$ since $p$ is prime to the order of $O_{K}^{\times}$. Hence for every $g\in\mathrm{Gal}(K(E[p])/K(p))$, there exists a unique $\tilde{g}\in\mathrm{Aut}_{K}(X)$ such that $t(g)$ and $\tilde{g}$ coincide in $\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)=\mathrm{Aut}(E[p])$. Since $t(g)\tilde{g}^{-1}$ has a prime-to-$p$ order and is contained in $\ker\left(\mathrm{Out}(\Pi_{1,1})\to\mathrm{Out}(\Pi_{1,1}^{\rm{ab}}/p)\right)$ which is pro-$p$, we conclude $t(g)=\tilde{g}$.
It shows that the image of the section $t$ coincides with a subgroup of $\mathrm{Aut}_{K}(X)$, which is contained in the center of $\rho_{X,p}(G_{K})$. In particular, the section $t$ induces the decomposition
$\rho_{X,p}(G_{K(p)})=\rho_{X,p}(G_{K(E[p])})\times\mathrm{imz}(t).$
@@
$\rho_{X,p}(G_{K(p)})=\mathrm{Gal}(\bar{K}^{\ker(\rho_{X,p})}/K(p))$
admits the required decomposition. ∎
@@
Recalling Sharifi’s result, one may wonder if this question has also an affirmative answer, assuming that Conjecture 2.8 holds and that a certain arithmetic condition on $p$ is satisfied. This is exactly what we prove in the following:
@@
    there are exactly two primes of $K(p^{\infty})$ above $p$, and
@@
Then the field $\Omega^{\ast}$ coincides with $\Omega$.
@@ 3.1. Two-variable filtration on free pro-$p$ group of rank two
Remark 3.1.
Let $\mathcal{L}$ be a free graded Lie algebra of rank two generated by $\{x_{\mathcal{L}},y_{\mathcal{L}}\}$. The filtration given in Definition 3.1 may be regarded as a group-theoretic analogue of the bi-grading on $\mathcal{L}$ induced by degrees with respect to $x_{\mathcal{L}}$ and $y_{\mathcal{L}}$.
@@
(3) We prove the assertion by induction on $|\boldsymbol{m}|+|\boldsymbol{n}|$. If $|\boldsymbol{m}|+|\boldsymbol{n}|=2$, then the assertion is clear. Let us assume that $|\boldsymbol{m}|+|\boldsymbol{n}|>2$. Since $\Pi(\boldsymbol{m})$ is normally generated by $[\Pi(\boldsymbol{m}^{\prime}),\Pi(\boldsymbol{m}^{\prime\prime})]$ with $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$, the assertion follows if there exist $\boldsymbol{n}^{\prime}$ and $\boldsymbol{n}^{\prime\prime}$ such that $\boldsymbol{m}^{\prime}\geq\boldsymbol{n}^{\prime}$, $\boldsymbol{m}^{\prime\prime}\geq\boldsymbol{n}^{\prime\prime}$ and $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$, by induction hypothesis. Such a pair $(\boldsymbol{n}^{\prime},\boldsymbol{n}^{\prime\prime})$ clearly exists unless $(m_{1},n_{1})$ or $(m_{2},n_{2})$ is equal to $(1,0)$. However, the assertion in this exceptional case also follows by using (2). ∎
@@
For $\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define graded quotients $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$ of $\Pi$ as
$\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(1,0))\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(0,1)).$
Note that $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$) is a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$ (resp. $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$)-module where the group $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$) acts by conjugation.
@@
(1) $\mathrm{gr}^{(1,0)}_{1}\Pi=\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module of rank one generated by $x$, cf. [Iha86, Theorem 2.2]. Similarly, $\mathrm{gr}^{(0,1)}_{2}\Pi$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$-module of rank one generated by $y$.
(2) $\mathrm{gr}^{(1,0)}_{2}\Pi=\Pi(1,0)/\Pi(\boldsymbol{1})$ is a free $\mathbb{Z}_{p}$-module of rank one generated by the image of $x$ on which $\Pi/\Pi(1,0)$ acts trivially. Similarly, $\mathrm{gr}^{(0,1)}_{1}\Pi$ is generated by $y$ on which $\Pi/\Pi(0,1)$ acts trivially. On the other hand, by Lemma 3.2(2), we have $\mathrm{gr}^{(m,0)}_{2}\Pi=\mathrm{gr}^{(0,m)}_{1}\Pi=0$ for every $m\geq 2$.

### v3

@@ 2.2. Elliptic analogues of Soulé characters
Moreover, the submodule $\Pi_{1,1}(m+3)\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ corresponds to $J^{m+1}$, where $J$ is the augmentation ideal of $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ [Iha86, (19) on page 67]. Hence the power series $T_{1}\alpha_{1,1}(\sigma)$ is contained in $J^{m+1}$, and we have $\alpha_{1,1}(\sigma)\in J^{m}$. This means that every coefficient of $\alpha_{1,1}(\sigma)$ of a monomial of total degree less than $m$ vanishes. By observing Theorem 2.6, it follows that $\kappa_{\boldsymbol{n}}(\sigma)$ vanishes for every $\boldsymbol{n}\in I$ such that $\lvert\boldsymbol{n}\rvert\leq m-2$ as desired. ∎
By Lemma 2.10 (2), we obtain a homomorphism
$\mathfrak{g}_{m}\xrightarrow{\oplus\kappa_{\boldsymbol{m}}}\bigoplus_{\boldsymbol{m}\in I,|\boldsymbol{m}|=m}\mathbb{Z}_{p}(\boldsymbol{m})$
for each even integer $m>2$. Since $\kappa_{\boldsymbol{m}}$ is originally defined on $F^{1}G_{K}$, this homomorphism factors though the $m$-th component $(\mathfrak{g}/[\mathfrak{g},\mathfrak{g}])_{m}$ of the abelianization of $\mathfrak{g}$. The following proposition tells us how to choose a basis of $\mathfrak{g}\otimes\mathbb{Q}_{p}$ in terms of elliptic Soulé characters:
Proposition 2.11.
Assume that the character $\kappa_{\boldsymbol{m}}$ is nontrivial for every $\boldsymbol{m}\in I$. Then, Conjecture 2.3 is equivalent to the following statement: For every even $m>2$, the homomorphism
$(\mathfrak{g}/[\mathfrak{g},\mathfrak{g}])_{m}\otimes\mathbb{Q}_{p}\xrightarrow{\oplus\kappa_{\boldsymbol{m}}}\bigoplus_{\boldsymbol{m}\in I,|\boldsymbol{m}|=m}\mathbb{Q}_{p}(\boldsymbol{m})$
is an isomorphism. Moreover, if we choose $\sigma_{\boldsymbol{m}}$ for each $\boldsymbol{m}\in I$ so that
  1. (a)
    $\sigma_{\boldsymbol{m}}$ is contained in the $\chi^{\boldsymbol{m}}$-isotypic component of $\mathfrak{g}_{\boldsymbol{\lvert m\rvert}}$, and
  2. (b)
    $\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})\neq 0$.
Then the elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ freely generate $\mathfrak{g}\otimes\mathbb{Q}_{p}$.
Proof.
Assume Conjecture 2.3. The first statement is a consequence of Lemma 2.10 (1). The second immediately follows from the first, since $\mathfrak{g}\otimes\mathbb{Q}_{p}$ is assumed to be free and the elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ generate $\mathfrak{g}$ (since they do at the level of the abelianization). ∎
2.3. Main result
We turn our attention to field-theoretic properties of the fixed field of the kernel of the pro-$p$ outer Galois representation. First, we study its ramification behavior as an extension over $K(E[p])$:
@@ 2.3. Main result
The field $\bar{K}^{\operatorname{Ker}(\rho_{X,p})}$ is a pro-$p$ extension of $K(E[p])$ unramified outside $p$.
@@
Since the kernel of the homomorphism $\operatorname{Out}(\Pi_{1,1})\to\operatorname{Aut}(\Pi_{1,1}^{\rm{ab}}/p)$ is pro-$p$, the field $\bar{K}^{\operatorname{Ker}(\rho_{X,p})}$ is a pro-$p$ extension of $K(E[p])$. To prove that the extension is unramified outside $p$, it suffices to show that $X\times_{\operatorname{Spec}(K)}\operatorname{Spec}(K(E[p]))$ or, equivalently, $E\times_{\operatorname{Spec}(K)}\operatorname{Spec}(K(E[p]))$ has good reduction everywhere outside $p$, by virtue of the specialization isomorphism [SGA71, X, Corollaire 3.9]. Using the Néron-Ogg-Shafarevich criterion, it suffices to show that every inertia subgroup $I$ associated with a prime of $K(E[p])$ lying outside $p$ acts trivially on $T_{p}(E)$. Since $E$ has complex multiplication, it has everywhere potentially good reduction. Hence the image of $I$ in $\operatorname{Aut}(T_{p}(E))$ is finite. Moreover, the image of $I$ is contained in $\operatorname{Ker}(\operatorname{GL}(T_{p}(E))\to\operatorname{GL}(E[p]))$, which is a torsion-free pro-$p$ group for every $p\geq 3$. Hence the image of $I$ is trivial as desired. ∎
Observe that the automorphism group $\operatorname{Aut}_{K}(X)$ defines a subgroup of $\operatorname{Out}(\Pi_{1,1})$. Since it is defined over $K$, it centralizes the image of $\rho_{X,p}$. This puts a strong constraint on the field $\bar{K}^{\operatorname{Ker}(\rho_{X,p})}$ as the following lemma shows:
@@
The field $\bar{K}^{\operatorname{Ker}(\rho_{X,p})}$ is a compositum of $K(E[p])$ and a subfield $\Omega^{\ast}\subset\Omega$ in such a way that $\rho_{X,p}(G_{K(p)})$ splits into the direct product
$\rho_{X,p}(G_{K(p)})=\operatorname{Gal}(\Omega/K(p))=\operatorname{Gal}(K(E[p])/K(p))\times\operatorname{Gal}(\Omega^{\ast}/K(p)).$
@@
$1\to\rho_{X,p}(G_{K(E[p])})\to\rho_{X,p}(G_{K(p)})\to\operatorname{Gal}(K(E[p])/K(p))\to 1.$
Since $\rho_{X,p}(G_{K(E[p])})$ is a pro-$p$ group and $\operatorname{Gal}(K(E[p])/K(p))$ is a finite prime-to-$p$ group (here we use $p\geq 5$), the exact sequence splits by the Schur-Zassenhaus theorem [Ser97, 5.9, Corollary 1]. Let us choose an arbitrary section
$t\colon\operatorname{Gal}(K(E[p])/K(p))\to\rho_{X,p}(G_{K(p)}).$
There is a natural homomorphism $O_{K}^{\times}=\operatorname{Aut}_{K}(X)\to\operatorname{Out}(\Pi_{1,1})$ by functoriality of étale fundamental groups, and the image of $\operatorname{Aut}_{K}(X)$ centralizes $\rho_{X,p}(G_{K})$. The $G_{K}$-action on $E[p]$ allows us to identify $\operatorname{Gal}(K(E[p])/K(p))$ with a subgroup of the image of $\operatorname{Aut}_{K}(X)$ in $\operatorname{Aut}_{O_{K}}(E[p])=(O_{K}/p)^{\times}$. Moreover, the automorphism group $\operatorname{Aut}_{K}(X)$ injects into $(O_{K}/p)^{\times}$ since $p$ is prime to the order of $O_{K}^{\times}$. Hence for every $g\in\operatorname{Gal}(K(E[p])/K(p))$, we can find a unique automorphism $\tilde{g}$ of $X$ such that $t(g)$ and $\tilde{g}$ induce the same element in $\operatorname{Out}(\Pi_{1,1}^{\rm{ab}}/p)=\operatorname{Aut}(E[p])$. Since $t(g)\tilde{g}^{-1}$ has a prime-to-$p$ order and is contained in the pro-$p$ group $\operatorname{Ker}\left(\operatorname{Out}(\Pi_{1,1})\to\operatorname{Out}(\Pi_{1,1}^{\rm{ab}}/p)\right)$, $t(g)$ must coincide with $\tilde{g}$. This argument shows that the image of $t$ is contained in (the image of) $\operatorname{Aut}_{K}(X)$. Therefore, $t$ induces the decomposition
$\rho_{X,p}(G_{K(p)})=\rho_{X,p}(G_{K(E[p])})\times\operatorname{Im}(t).$
@@
$\rho_{X,p}(G_{K(p)})=\operatorname{Gal}(\bar{K}^{\operatorname{Ker}(\rho_{X,p})}/K(p))=\operatorname{Gal}(K(E[p])/K(p))\times\operatorname{Gal}(\Omega^{\ast}/K(p))$
is the required decomposition. ∎
@@
In analogy with Sharifi’s result, one may wonder if the question has an affirmative answer, assuming Conjecture 2.3 and a certain arithmetic assumption on $p$ is satisfied. This is exactly the main result of the present paper:
@@
    there are exactly two primes of $K(p^{2})$ above $p$, and
@@
Then $\Omega^{\ast}$ coincides with $\Omega$. Consequently, we have $\bar{K}^{\operatorname{Ker}(\rho_{X,p})}=K(E[p])\cdot\Omega$.
The proof of Theorem 2.14 will be given at the end of Section 4.
Remark 2.14.1.
We keep the same notation as in Theorem 2.14, and write $p=\pi\bar{\pi}$ for a prime element $\pi\in K$ and its conjugate $\bar{\pi}$. Then the following statements are equivalent:
  1. (1)
    There are exactly two primes of $K(p^{2})$ above $p$.
  2. (2)
    $\operatorname{Gal}(K(p^{2})/K)$ coincides with its decomposition subgroup at $\mathfrak{p}$.
  3. (3)
    The image of $\pi$ in $(O_{K}/\bar{\mathfrak{p}}^{2})^{\times}/O_{K}^{\times}$ is a generator of this cyclic group.
  4. (4)
    $\operatorname{Gal}(K(p^{\infty})/K)$ coincides with its decomposition subgroup at $\mathfrak{p}$.
  5. (5)
    The image of $\pi$ in $O_{K_{\bar{\mathfrak{p}}}}^{\times}/O_{K}^{\times}$ is a (topological) generator of this procyclic group.
Observe that we do not use the assumption that $p$ splits in $K$ to prove Lemma 2.13 (though we use the fact that $p$ is prime to the order of $O_{K}^{\times}$). We expect an analogue of Anderson-Ihara’s question to have an affirmative answer even if $p$ does not split. For example, we can show the following unconditional result, whose proof relies on the case of genus zero (this result is not used in the rest of the paper):
Proposition 2.15.
Suppose
$K\in\{\mathbb{Q}(\sqrt{-7}),\mathbb{Q}(\sqrt{-11}),\mathbb{Q}(\sqrt{-19}),\mathbb{Q}(\sqrt{-43}),\mathbb{Q}(\sqrt{-167})\}$
and let $p_{K}$ be a unique rational prime which ramifies in $K$. Then we have $\bar{K}^{\operatorname{Ker}(\rho_{X,p_{K}})}=K(E[p_{K}])\cdot\Omega$.
Proof.
By our assumption, we have
$p_{K}\in\{7,11,19,43,163\}$
and hence $p_{K}$ is odd and regular. Moreover, it is easy to observe that the mod-$p$ ray class field $K(p_{K})$ is a $p_{K}$-extension of $\mathbb{Q}(\mu_{p_{K}})=K(\mu_{p_{K}})$ unramified outside $p_{K}$, since it contains $K(\mu_{p_{K}})$ and $K(p):K(\mu_{p})]=p$. On the other hand, by ‘‘Oda’s prediction’’ [Tak12, Theorem 3.6] established by Takao, it follows that
$\bar{\mathbb{Q}}^{\operatorname{Ker}(\rho_{\mathbb{P}^{1}_{\mathbb{Q}}\setminus\{0,1,\infty\},p_{K}})}\subset\bar{K}^{\operatorname{Ker}(\rho_{X,p_{K}})}.$
However, the former field coincides with $\Omega^{\mathrm{cyc}}$ by Sharifi’s theorem [Sha02, Theorem 1.1], together with the resolution of the Deligne-Ihara conjecture by Hain-Matsumoto [HM03] and Brown [Bro12]. Since $\Omega^{\mathrm{cyc}}$ is also the maximal pro-$p_{K}$ extension of $K(p_{K})$ unramified outside $p_{K}$, it coincides with $\Omega$. This concludes the proof since the proof of Lemma 2.13 works even when $p\geq 5$ ramifies in $K$. ∎
Remark 2.15.1.
Proposition 2.15 even holds for $K=\mathbb{Q}(\sqrt{-3})$ (and hence $p_{K}=3$); In fact, since $E$ is then isomorphic to the Fermat curve of degree 33 over $K(E[3])$, we may assume that $E$ is the Fermat curve of degree $3$. Then it is easy to observe $K=K(3)=K(E[3])$. B Lemma 2.12 and a similar argument as that of Proposition 2.15, we have $\Omega^{\mathrm{cyc}}=\Omega\subset\bar{K}^{\operatorname{Ker}(\rho_{X,3})}\subset\Omega$ as desired.
@@ 3.1. Two-variable filtration on free pro-$p$ group of rank two
(3) We prove the assertion by induction on $|\boldsymbol{m}|+|\boldsymbol{n}|$. If $|\boldsymbol{m}|+|\boldsymbol{n}|=2$, then the assertion is clear. Let us assume that $|\boldsymbol{m}|+|\boldsymbol{n}|>2$. Since $\Pi(\boldsymbol{m})$ is normally generated by $\Pi(\boldsymbol{m}^{\prime}),\Pi(\boldsymbol{m}^{\prime\prime})]$ with $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$, the assertion follows if there exist $\boldsymbol{n}^{\prime}$ and $\boldsymbol{n}^{\prime\prime}$ such that $\boldsymbol{m}^{\prime}\geq\boldsymbol{n}^{\prime}$, $\boldsymbol{m}^{\prime\prime}\geq\boldsymbol{n}^{\prime\prime}$ and $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$, by induction hypothesis. Such a pair $(\boldsymbol{n}^{\prime},\boldsymbol{n}^{\prime\prime})$ clearly exists unless $(m_{1},n_{1})$ or $(m_{2},n_{2})$ is equal to $(1,0)$. However, the assertion in this exceptional case also follows by using (2). ∎
@@
For $\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define graded quotients $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ and $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Pi$ of $\Pi$ as
$\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(1,0))\quad\text{and}\quad\operatorname{Gr}^{\boldsymbol{m}}_{2}\Pi\coloneqq\Pi(\boldsymbol{m})/\Pi(\boldsymbol{m}+(0,1)).$
Note that $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Pi$) is a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$ (resp. $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$)-module where the group $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$) acts by conjugation.
@@
(1) $\operatorname{Gr}^{(1,0)}_{1}\Pi=\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module of rank one generated by $x$, cf. [Iha86, Theorem 2.2]. Similarly, $\operatorname{Gr}^{(0,1)}_{2}\Pi$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(0,1)]]$-module of rank one generated by $y$.
(2) $\operatorname{Gr}^{(1,0)}_{2}\Pi=\Pi(1,0)/\Pi(\boldsymbol{1})$ is a free $\mathbb{Z}_{p}$-module of rank one generated by the image of $x$ on which $\Pi/\Pi(1,0)$ acts trivially. Similarly, $\operatorname{Gr}^{(0,1)}_{1}\Pi$ is generated by $y$ on which $\Pi/\Pi(0,1)$ acts trivially. On the other hand, by Lemma 3.3(2), we have $\operatorname{Gr}^{(m,0)}_{2}\Pi=\operatorname{Gr}^{(0,m)}_{1}\Pi=0$ for every $m\geq 2$.